OpenStudy (anonymous):

Factor the following expression.. 198q3r2 – 184q2r2 + 18qr2

OpenStudy (anonymous):

what can you factor out of ALL of them?

OpenStudy (anonymous):

uhmm... \[r ^{2}\]

OpenStudy (johnweldon1993):

that's one....what about the q's does each have at least 1 q?

OpenStudy (anonymous):

what does 198, - 184, and 18 have in common?

OpenStudy (anonymous):

Yes, that's true annnnd I'm not sure.... they all can be divided by two... orrrrr well idk

OpenStudy (johnweldon1993):

okay so we have r² and q so far 198/2 = 99 184/2 = 92 18/2 = 9 they cant be simplified any farther so yes 2qr²(99q² - 92q + 9) so how can we factor this?

OpenStudy (anonymous):

I have no idea